nevewpaatenp 202353 1888us6 id118570 1&content id 2487904 18return content-18step Question 5 0 out of 5 points Problem 3 Part show with solving please
Answer #1

Effective annual rate = (Increase in value)*100/initial value

= (13.14 - 6.90)*100/6.90

= 90.43%

For quarterly compounding use (1 + Effective annual rate)^(1/4) - 1

= (1+90.43%)^(1/4)-1

= 17.472%

For monthly compounding use (1 + Effective annual rate)^(1/12) - 1

= (1+90.43%)^(1/12) -1

= 5.514%